OpenStudy (anonymous):

$$if \quad u \quad = \left(\frac{\left(\arctan \left(xy\right)\right)}{\left(\sqrt{1+x^2+y^2}\right)}\right)$$ prove $$\frac{\partial{u}}{\partial{x}}=\frac{y}{(1+x^2)(\sqrt{1+x^2+y^2}}$$

OpenStudy (anonymous):

i got $$\frac{\frac{y}{y^2x^2+1}\sqrt{1+x^2+y^2}-\frac{x}{\sqrt{1+x^2+y^2}}\arctan \left(xy\right)}{\left(\sqrt{1+x^2+y^2}\right)^2}$$ using quotient rule

OpenStudy (anonymous):

college stuff :{

zepdrix (zepdrix):

It looks like what `you got` is \(\rm u\) not \(\rm u_x\). Notice that \(\rm u\) is the first derivative with respect to x, yes?

zepdrix (zepdrix):

first partial*

zepdrix (zepdrix):

Hmm, second derivative is going to be a bit of work :[ I gotta head to class unfortunately.

OpenStudy (anonymous):

sorry!! I've entered differential operator wrongly. the question is corrected now.